# Equilibrium of Charges and Differential Equations Solved by Polynomials II

### v1

@@ 7 Conclusions and Open Problems
In the third-order case, one has to deal with Pfaffians, rather than Wronskians [23, 15]. “Pfaffian analog” of the Crum theorem exists for a sequence of Darboux transformations associated with the set of distinct eigenvalues [23]. However, it is unclear how to deal with its zero-level degeneration, because, in difference from the second-order case, the eigenfunctions must satisfy certain boundary conditions [5, 23, 4].
Recently, a Pfaffian representation of polynomial $\tau$-functions of the Sawada-Kotera hierarchy was found by Kac and Van de Leur in [15] via free-fermion approach.
@@
It was shown by Fordy and Gibbons [12] that solutions $u$ of the Sawada-Kotera and solutions $v$ of the Kaup-Kupershmidt hierarchy are related by the Miura transormation
$u=w^{\prime}+w^{2},\quad v=w^{\prime}-\frac{1}{2}w^{2}.$
@@
Finally, we would like to outline several generalizations of systems considered in this paper: One of generalizations extends (1) to the case of translating configurations or equilibria of charges in the homogeneous electric field $k$:
@@
When $\Lambda=1$, $\Phi(z,k)$ is the “rational” Baker-Akhiezer function of the KdV hierarchy. It can be obtained by application of sequence of the $\eta=0$ Darboux transformations (13, 14) to the eigenfunction $\psi=e^{kz}$ of the free Schrodinger operator corresponding to the eigenvalue $\lambda=k^{2}$ (for review see e.g. [21] and references therein).
@@
When $\Lambda=1$, this generalization is associated with the soliton solutions of the KdV hierarchy and “trigonometric” Baker-Akhieser function $\Phi=q(z,k)/p(z,k)e^{kz}$, where now $q$ and $p$ are trigonometric polynomials in $z$ ([16, 20, 13], for review see e.g. [21]). Trigonometric Baker-Akhiezer function is obtained from the eigenfunction $e^{kz}$ (with eigenvalue $\lambda=k^{2}$) of the free Schrodinger operator by iterations of the Darboux transformations: First with $\eta=k_{1}^{2}$, then transformation with $\eta=k_{2}^{2}$, etc. Here $k_{i}$ are integers, such that $0<k_{1}<k_{2}<\dots$.
When $\Lambda=2$, vortex streets configurations cannot be generated by Darboux transformations, even in the non-translating case $k=0$ only short terminating sequences can be produced by the zero-level transformations. Indeed, similarly to the second order case, operators $L$ with periodic coefficients can be obtained by sequence of Darboux transformations from $\partial^{3}$ only when transformations are associated with a sequence of distinct, nonzero $\eta$. Here equations
@@
where $u_{i}$ is periodic and $k_{i}\not=0$, cannot be reduced to the Schrodinger equations. As a consequence, $\varkappa_{i}$, do not correspond to equilibrium configurations . It would be interesting to explore the $\Lambda=2$ vortex street configurations using alternative methods.

### v2

@@ 7 Conclusions and Open Problems
In the third-order case, one has to deal with Pfaffians, rather than Wronskians [16, 24]. “Pfaffian analog” of the Crum theorem exists for a sequence of Darboux transformations associated with the set of distinct seed eigenvalues [24]. However, it is unclear how to deal with its zero-level degeneration, because, in difference from the second-order case, the eigenfunctions must satisfy certain boundary conditions [4, 5, 24].
Recently, a Pfaffian representation of polynomial $\tau$-functions of the Sawada-Kotera hierarchy was found by Kac and Van de Leur in [16] via free-fermion approach (for more details on Sawada-Kotera and Kaup-Kupershmidt hierarchies, see e.g. [4, 13, 15, 16] and references therein).
@@
It was shown by Fordy and Gibbons [13] that solutions $u$ of the Sawada-Kotera and solutions $v$ of the Kaup-Kupershmidt hierarchy [^7] The Lax operator of the Kaup-Kupersmidt hierarchy equals $L=\partial^{3}+v\partial+\frac{1}{2}v^{\prime}$ are related by the Miura transormation
$u=w^{\prime}+w^{2},\quad v=2w^{\prime}-w^{2}.$
@@
Finally, we would like to outline several generalizations of systems considered in this paper: One of generalizations extends (1) to the case of configurations translating with constant velocity or equilibria of charges in homogeneous background electric field $k$:
@@
When $\Lambda=1$, $\Phi(z,k)$ is the “rational” Baker-Akhiezer function of the KdV hierarchy. It can be obtained by application of sequence of the $\eta=0$ Darboux transformations (14, 15) to the eigenfunction $\psi=e^{kz}$ of the free Schrodinger operator corresponding to the eigenvalue $\lambda=k^{2}$. In this way one gets all possible translating configurations for $Q_{i}\in\{-1,1\}$ (see e.g. [22]).
@@
When $\Lambda=1$, this generalization is associated with the soliton solutions of the KdV hierarchy and “trigonometric” Baker-Akhieser function $\Phi=q(z,k)/p(z,k)e^{kz}$, where now $q$ and $p$ are trigonometric polynomials in $z$ ([14, 17, 21], for review see e.g. [22]). Trigonometric Baker-Akhiezer function is obtained from the eigenfunction $e^{kz}$ (with eigenvalue $\lambda=k^{2}$) of the free Schrodinger operator $\partial^{2}$ by iterations of the Darboux transformations: First with $\eta=k_{1}^{2}$, then transformation with $\eta=k_{2}^{2}$, etc. Here $k_{i}$ are integers, such that $0<k_{1}<k_{2}<\dots$.
When $\Lambda=2$, translating vortex streets configurations cannot be generated by Darboux transformations: Even in the non-translating case $k=0$, only short terminating sequences can be produced by the zero-level transformations [22]. Indeed, similarly to the second order case, operators $L$ with non-zero periodic coefficients can be obtained by sequence of Darboux transformations from $\partial^{3}$ only when transformations are associated with a sequence of distinct, nonzero $\eta$. Here equations
@@
where $u_{i}$ is periodic and $k_{i}\not=0$, cannot be reduced to the Schrodinger equations. As a consequence, $\varkappa_{i}$ do not correspond to equilibrium configurations. It would be interesting to explore the $\Lambda=2$ vortex street configurations using alternative methods.